OpenStudy (anonymous):

N^2 = 6n+c , Solve for c to make perfect square

OpenStudy (anonymous):

Cute

OpenStudy (anonymous):

Why is there an \(=\) sign?

OpenStudy (anonymous):

Should it be \(+\) or \(-\)?

OpenStudy (anonymous):

Im not sure.. Whats cute about this..

OpenStudy (anonymous):

Is it really an \(=\)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Then I think the first thing you have to do is move it all to one side.

OpenStudy (anonymous):

Its a quadratic equation..

OpenStudy (anonymous):

\[ n^2-6n-c=0 \]

OpenStudy (anonymous):

In general, we would observe that \[ (n+a)^2 = n^2+an+an+a^2 = n^2+2an+a^2 \]

OpenStudy (anonymous):

If we set them equal, then: \[ n^2+2an+a^2=n^2-6n-c \]And if we set the coefficients equal, we get: \[ 2a=-6\\ a^2=-c \]

OpenStudy (anonymous):

When we solve the first equation\[ 2a=-6\implies a=-6/2=-3 \]We can put it into the second equation: \[ (-3)^2 = -c\implies 9=-c\implies c=-9 \]

OpenStudy (anonymous):

So we would say: \[ \big (n-3\big)^2 = n^2-6n+9 \]And that \(c=-9\).

OpenStudy (anonymous):

I think this question is difficult for no reason, but thank you.

OpenStudy (anonymous):

Usually they have you memorize some formula

OpenStudy (anonymous):

Like \(b/2\) or something